OpenStudy (caylie00):

I need help will fan and other . Simplify the radical expression √5over√8x

jimthompson5910 (jim_thompson5910):

So the original expression is this? \[\Large \frac{\sqrt{5}}{\sqrt{8x}}\]

OpenStudy (caylie00):

yes

jimthompson5910 (jim_thompson5910):

Multiply top and bottom by \(\Large {\color{red}{\sqrt{8x}}}\) to rationalize the denominator \[\Large \frac{\sqrt{5}}{\sqrt{8x}} = \frac{\sqrt{5}{\color{red}{*\sqrt{8x}}}}{\sqrt{8x}{\color{red}{*\sqrt{8x}}}}\] Do you see how to finish up?

OpenStudy (caylie00):

um i think so..

jimthompson5910 (jim_thompson5910):

What would \[\Large \sqrt{8x}*\sqrt{8x}\] simplify to?